Can u help me with this word problem.... the minute hand of a clock is 8in long and moves from12 to 2 o'clock. How far does the tip of the minute hand move? express the answer in term of pie and then round to two deciaml places....
Answer by stanbon(75887)

the minute hand of a clock is 8in long and moves from12 to 2 o'clock. How far does the tip of the minute hand move? express the answer in term of pie and then round to two deciaml places....
---------
Circumference = 2(pi)r
Your Problem:
circumference = 16(pi) inches
-------------
From 12:00 PM to 2:00PM the minute hand travels twice around the clock
Distance traveled is 2*16(pi) = 32(pi) = 100.53 inches
================
Cheers,
Stan H.